OpenStudy (anonymous):

What is 3 7/10 - 2 1/2

OpenStudy (anonymous):

ive done this problem already - 1 1/5

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

i think the asnwer is 1 and 1/5 (NOT negative...)

jhonyy9 (jhonyy9):

37/10 - 5/2 = (37-25)/10 = 12/10 = 6/5

OpenStudy (anonymous):

yep moyo i meant ' -' as a hyphen - the answer is 1 1/5 or 6/5